Light traveling through Air encounters a second medium which slows the light to 2.7×10 to the power of eight minutes per second what is the index of the second medium
1
Expert's answer
2019-06-17T12:54:27-0400

the index of the second medium is given by formula


"n= \frac { c } {v} (1)"

where c=3×108 m/s, v=2.7×108 m/s

Using (1) we get: n=1.11


Answer:

1.11